# NUTSHELL: A Dataset for Abstract Generation from Scientific Talks

### v1

@@ 1 Introduction
Scientific summarization has been widely studied in natural language processing, including summarizing entire articles (Collins et al., 2017; Liu et al., 2024), particularly in the medical domain (Kedzie et al., 2018; Cohan et al., 2018; Gupta et al., 2021), generating abstracts from citations (Yasunaga et al., 2019; Zanzotto et al., 2020), summarizing specific paper sections (Takeshita et al., 2024), and leveraging knowledge graphs for abstract generation (Koncel-Kedziorski et al., 2019).
@@
Table 1: Dataset statistics. The number of words is obtained by splitting the abstract at whitespaces.
@@ 2 The NUTSHELL Dataset
In this section, we introduce the new NUTSHELL resource. We chose to build our corpus upon the the ACL Anthology [^2] https://aclanthology.org since it provides a rich collection of multimodal resources (talks and abstracts) and open-access licensing. Starting from 2017, a significant number of papers published in the main *ACL conferences (ACL, EMNLP, and NAACL) include a video of the presentation, all released under the Creative Commons Attribution 4.0 license. This makes ACL an ideal resource for building a multimodal dataset for the SAG task.
@@ 2.1 Are paper abstracts “good” talk summaries?
Before creating the corpus, we establish the validity of our data by investigating whether abstracts represent a good summary of the associated talk. To this aim, we conduct a qualitative check on a data sample of 30 talk-abstract pairs from the ACL anthology. We involve a total of 5 annotators, who are all domain experts and thus familiar with scientific material. [^3] Annotators include the paper authors and their colleagues, whose work will be acknowledged upon acceptance. To verify Inter-Annotator Agreement (IAA), a double annotation by different experts was carried out on 15 pairs.
Since we are interested in understanding whether paper abstracts are informative enough to represent a good summary of the talk, we asked evaluators to annotate: (1) Whether the information in the abstract is all uttered by the presenter; (2) The span of information present in the abstract that was not contained in the talk, if any; (3) Whether they think that the abstract summarizes all important information presented in the talk. The human evaluation template is provided in Fig. 2 of App. A.
The results indicate that $70.0\%$ of the abstracts are considered good summaries by annotators as they contain important information about the talk. However, $63.3\%$ of the abstracts contain information that is not present in the talk. For this reason, we analyzed the annotated spans of the missing information. We observed that this phenomenon is mainly due to missing dataset, model, and shared task (e.g., evaluation campaigns) names or URLs (e.g., link to the resource or model being released), which are typically not spelled by presenters. [^4] This issue could be overcome by exploiting the videos, as this information is typically shown in the slides. While out of scope for SAG, NUTSHELL includes the videos, making it a useful resource also for more complex multimodal tasks. Despite this drawback, the evaluation of automatic models against the same ground truth abstract can be considered fair, as models are equally penalized by this category of missing information. Moreover, establishing a unique ground truth for summarization tasks is still an open research question (Zhang et al., 2024) as humans often produce very different summaries. Both, questions (1) and (3) have an inter-annotator agreement of $\kappa=0.466$, indicating moderate agreement (Landis and Koch, 1977), which is satisfactory given the subjective nature of evaluating summaries. Therefore, the manual evaluation revealed the feasibility of the SAG tasks and the validity of our resource.
@@ 2.2 Collection and Dataset Statistics
We collected talks from 16 ACL Anthology events: 6 ACL, 6 EMNLP, and 4 NAACL, including workshops. For each paper (both long and short format), we extracted the video and the associated abstract already available on the paper website. We exclude papers with invalid URLs, videos without audio, or abstracts missing from the paper page.
Lastly, we split the dataset into training (years 2017 to 2021), dev (ACL 2022), and test (EMNLP/NAACL 2022). These splits reflect a realistic evaluation setup, where models are trained on past data and tested on the most recent, unseen examples. In total, the corpus contains 1,172 hours of audio content corresponding to 6,316 different presentations (full statistics are reported in Table 1).
@@ 3.1.1 Models
To establish baselines for the SAG task, we analyze the performance of four models described as follows. Prompts, model, generation, and additional training details are provided in App. B.
@@
Qwen2-Audio-7B-Instruct.
End2End Zero-Shot.
A SpeechLLM composed of HuBERT (Hsu et al., 2021) as speech encoder, meta-llama/Llama-3.1-8B-Instruct as LLM, and a QFormer (Li et al., 2023) as adapter. The SpeechLMM is built to handle long audio inputs (App. B) and obtained by training only the adapter in two steps: (a) contrastive pretraining (Züfle and Niehues, 2024) to align the LLM representations for the speech and text modalities using MuST-C (Di Gangi et al., 2019) and Gigaspeech (Chen et al., 2021), and (b) fine-tuning on instruction-following tasks, including ASR, speech translation, and spoken question answering using MuST-C and Spoken-SQuAD (Lee et al., 2018). Therefore, the model is not trained or fine-tuned on NUTSHELL and operates in zero-shot for the SAG task.
End2End Finetuned.
@@ 3.1.2 Evaluation
We use standard (text) summarization metrics: ROUGE (Lin, 2004) – a text similarity metric that has been widely adopted for LM evaluation (Grusky, 2023) that focuses on n-gram overlap between the hypothesis and reference –, and BERTScore (Zhang et al., 2020) – a neural-based metric that measures the pairwise similarity of contextualized token embeddings between the summary and its reference. Also, we rely on LLM-as-a-judge (Shen et al., 2023; Zheng et al., 2024) where the LLM [^6] We use Llama-3.1-8B-Instruct (Dubey et al., 2024) as the judge using the prompts reported in Fig. 1 in Section C.2. is prompted to assign a score to each output, using the reference abstract as context (Score with Expl.). The score is based on four criteria: (1) relevance, (2) coherence, (3) conciseness, and (4) factual accuracy.777(1) Does the predicted abstract capture the main points of the gold abstract?, (2) Is the predicted abstract logically organized and easy to follow?, (3) Is the predicted abstract free from unnecessary details?, (4) Are the claims in the predicted abstract consistent with the gold abstract? We also report results where the LLM judge provides a single score without explanations (Plain Score), as well as results where it ranks the given abstracts instead of scoring them individually (Avg. Rank).
@@
For the human evaluation, 9 annotators – all experts in the field – were provided with the generated abstracts and the ground truth abstract. We use the same randomly sampled 30 test set examples as in Section 2.1 and validate their representativeness, which is discussed in App. D. Each sample is evaluated by three annotators. They follow the same criteria as the LLM evaluation but rank models instead of assigning scores. Detailed instructions are in App. D. As the End2End Zero-Shot model performance was comparable to that of Qwen2-Audio – also being a zero-shot model – and given that Qwen2-Audio is an established SpeechLLM with a distinct architecture, we exclude the End2End Zero-Shot from this analysis.
@@ 3.2 Results
For a more granular analysis, Table 3 in Section C.2 provides results for the LLM-based metrics. Given that all models except Qwen2-Audio rely on Llama3.1-8B-Instruct, one might question whether the Llama-based judge could introduce bias in favor of these models. To address this, we perform additional evaluations using Qwen/Qwen2-7B (Yang et al., 2024) as the judge (Table 4 in Section C.2), which confirm the same ranking, eliminating any concerns about evaluator bias.
@@ Appendix A Human Evaluation: Are abstracts good summaries of the talk?
We want to analyze whether the paper abstracts can serve as good abstracts for the *ACL talks. Therefore, we conduct a human evaluation, sampling 30 examples from our dataset. Detailed instructions for the human annotators can be found in Fig. 2.
Appendix B Baseline Details
@@
Appendix C Evaluation Details
@@
C.1 ROUGE and BERT Score
@@ C.2 LLM as a judge
Appendix D Human Evaluation for Ranking Model Outputs
We evaluate the models using ROUGE (Lin, 2004), BERTScore (Zhang et al., 2020), and LLM-as-a-judge. However, the first two metrics may not fully capture information overlap (Deutsch and Roth, 2021), while LLM-as-a-judge is sensitive to different prompts (Thakur et al., 2024) and struggles with distinguishing similar candidates (Shen et al., 2023). To address these limitations, we complement our evaluation with human analysis. Specifically, we ask 9 domain experts to rank model outputs relative to the reference abstract, with each example annotated by three independent annotators. The annotation instructions are provided in Fig. 3.

### v2

@@ 1 Introduction
Scientific summarization has been widely studied in natural language processing, including summarizing entire articles (Collins et al., 2017; Mao et al., 2022; Liu et al., 2024), particularly in the medical domain (Kedzie et al., 2018; Cohan et al., 2018; Gupta et al., 2021), generating abstracts from citations (Yasunaga et al., 2019; Zanzotto et al., 2020), summarizing specific paper sections (Takeshita et al., 2024), and leveraging knowledge graphs for abstract generation (Koncel-Kedziorski et al., 2019).
Figure 1: NUTSHELL, a dataset for Speech-to-Abstract Generation (SAG) from scientific talks.
@@
Table 1: Dataset statistics for NUTSHELL. The number of words is obtained by splitting the abstract at whitespaces.
@@ 2 The NUTSHELL Dataset
In this section, we introduce the new NUTSHELL resource. We chose to build our corpus upon the the ACL Anthology [^2] https://aclanthology.org since it provides a rich collection of multimodal resources (talks and abstracts) and open-access licensing. Starting from 2017, a significant number of papers published in the main *ACL conferences (ACL, EMNLP, and NAACL) include a video of the presentation, all released under the Creative Commons Attribution 4.0 license. This makes *ACL an ideal resource for building a multimodal dataset for the SAG task.
@@ 2.1 Are paper abstracts “good” talk summaries?
Before creating the corpus, we establish the validity of our data by investigating whether abstracts represent a good summary of the associated talk. To this aim, we conduct a qualitative check on a data sample of 30 talk-abstract pairs from the ACL Anthology. We involve a total of 5 annotators, who are all domain experts and thus familiar with scientific material. [^3] Annotators include the paper authors and their colleagues. To verify Inter-Annotator Agreement (IAA), a double annotation by different experts was carried out on 15 pairs.
Since we are interested in understanding whether paper abstracts are informative enough to represent a good summary of the talk, we asked evaluators to annotate: (1) Whether the information in the abstract is all uttered by the presenter in the talk; (2) The span of information present in the abstract that was not contained in the talk, if any; (3) Whether the abstract summarizes all important information presented in the talk. The human evaluation procedure, including the annotation template, is described in App. A.
The results indicate that $70.0\%$ of the abstracts are considered good summaries by annotators as they contain important information about the talk. However, $63.3\%$ of the abstracts also contain information not explicitly present in the talk itself. To better understand this, we conducted a qualitative analysis of the annotated spans corresponding to this missing information. We found that these spans typically involved dataset names, model names, shared task references (e.g., evaluation campaigns), or URLs (e.g., link to the resource or model being released). Notably, these elements are often displayed on slides but not explicitly verbalized by presenters. [^4] This issue could be overcome by exploiting the videos, as this information is typically shown in the slides. While out of scope for SAG, NUTSHELL includes the videos, making it a useful resource also for more complex multimodal tasks.
Despite this issue, the evaluation of automatic models against the same ground truth abstract can be considered fair, as models are equally penalized by this category of missing information. Moreover, it is worth noting that establishing a single ground truth for summarization tasks is still an open challenge (Zhang et al., 2024), given the inherent variability in human-produced summaries.
Both, questions (1) and (3) have an inter-annotator agreement of $\kappa=0.466$, indicating moderate agreement (Landis and Koch, 1977), which can be regarded as acceptable given the subjective nature of evaluating summaries. While criterion (3) naturally involves subjective judgments about information importance, the lower agreement on criterion (1) can also be attributed to borderline cases, where small phrasing differences were sometimes overlooked by individual annotators. Such subtleties led to occasional discrepancies in annotator decisions, but were manually reviewed.
In summary, the manual evaluation confirmed both the feasibility of the SAG tasks and, despite the noted challenges, the overall reliability and usefulness of our resource.
@@ 2.2 Collection and Dataset Statistics
We collected talks from 16 ACL Anthology events: 6 ACL, 6 EMNLP, and 4 NAACL, including workshops, shared tasks and industry tracks. For each paper (both long and short format), we extracted the video and the associated abstract already available on the paper website. We exclude papers with invalid URLs, videos without audio, or abstracts missing from the paper page. Additional details on the data collection can be found in App. B.
Lastly, we split the dataset into training (years 2017 to 2021), dev (ACL 2022), and test (EMNLP/NAACL 2022). These splits reflect a realistic evaluation setup, where models are trained on past data and tested on the most recent, unseen examples. In total, the corpus contains 1,172 hours of audio content corresponding to 6,316 different presentations. Full statistics are reported in Table 1.
@@ 3.1.1 Models
To establish baselines for the SAG task, we analyze the performance of four models described as follows. Prompts, model, generation, and additional training details are provided in App. C.
@@
Qwen2-Audio-7B-Instruct.
End2End Zero-Shot.
A SpeechLLM composed of HuBERT (Hsu et al., 2021) as speech encoder, meta-llama/Llama-3.1-8B-Instruct as LLM, and a QFormer (Li et al., 2023) as adapter. The SpeechLMM is built to handle long audio inputs (App. C) and obtained by training only the adapter in two steps: (a) contrastive pretraining (Züfle and Niehues, 2024) to align the LLM representations for the speech and text modalities using MuST-C (Di Gangi et al., 2019) and Gigaspeech (Chen et al., 2021), and (b) fine-tuning on instruction-following tasks, including ASR, speech translation, and spoken question answering using MuST-C and Spoken-SQuAD (Lee et al., 2018). Therefore, the model is not trained or fine-tuned on NUTSHELL and operates in zero-shot for the SAG task.
End2End Finetuned.
@@ 3.1.2 Evaluation
We use standard (text) summarization metrics: ROUGE (Lin, 2004) – a text similarity metric that has been widely adopted for LM evaluation (Grusky, 2023) that focuses on n-gram overlap between the hypothesis and reference –, and BERTScore (Zhang et al., 2020) – a neural-based metric that measures the pairwise similarity of contextualized token embeddings between the summary and its reference. Also, we rely on LLM-as-a-judge (Shen et al., 2023; Zheng et al., 2024) where the LLM [^6] We use Llama-3.1-8B-Instruct (Dubey et al., 2024) as the judge using the prompts reported in Fig. 2 in Section D.2. is prompted to assign a score to each output, using the reference abstract as context (Score with Expl.). The score is based on four criteria: (1) relevance, (2) coherence, (3) conciseness, and (4) factual accuracy.777(1) Does the predicted abstract capture the main points of the gold abstract?, (2) Is the predicted abstract logically organized and easy to follow?, (3) Is the predicted abstract free from unnecessary details?, (4) Are the claims in the predicted abstract consistent with the gold abstract? We also report results where the LLM judge provides a single score without explanations (Plain Score), as well as results where it ranks the given abstracts instead of scoring them individually (Avg. Rank).
@@
For the human evaluation, nine annotators – all experts in the field – were provided with the generated abstracts and the ground truth abstract. We use the same randomly sampled 30 test set examples as in Section 2.1 and validate their representativeness, which is discussed in App. E. Each sample is evaluated by three annotators. They follow the same criteria as the LLM evaluation but rank models instead of assigning scores. Detailed instructions are in App. E. As the End2End Zero-Shot model performance was comparable to that of Qwen2-Audio – also being a zero-shot model – and given that Qwen2-Audio is an established SpeechLLM with a distinct architecture, we exclude the End2End Zero-Shot from this analysis.
@@ 3.2 Results
For a more granular analysis, Table 3 in Section D.2 provides results for the LLM-based metrics. Given that all models except Qwen2-Audio rely on Llama3.1-8B-Instruct, one might question whether the Llama-based judge could introduce bias in favor of these models. To address this, we perform additional evaluations using Qwen/Qwen2-7B (Yang et al., 2024) as the judge (Table 4 in Section D.2), which confirm the same ranking, eliminating any concerns about evaluator bias.
@@ 5 Limitations
Acknowledgments
This work has received funding from the European Union’s Horizon research and innovation programme under grant agreement No 101135798, project Meetween (My Personal AI Mediator for Virtual MEETtings BetWEEN People). We gratefully acknowledge Poland’s high-performance Infrastructure PLGrid ACC Cyfronet AGH for providing computer facilities. Beatrice Savoldi and Marco Gaido are supported by the PNRR project FAIR - Future AI Research (PE00000013), under the NRRP MUR program funded by the NextGenerationEU.
We thank Leonard Bärmann, Béni Egressy, Mia Fuß, Lukas Hilgert, and Danni Liu for their contributions to the data annotation process.
@@ Appendix A Human Evaluation: Are abstracts good summaries of the talk?
We aim to assess whether paper abstracts can serve as effective abstracts for *ACL talks. To this end, we conducted a human evaluation by randomly sampling 30 examples from our dataset. The annotation team consisted of five individuals (four women and one man), including the paper authors and their colleagues. All annotators were already familiar with the NLP domain, scientific presentation and writing, and the task itself. They are experts in Natural Language Processing, holding at least a master’s degree in NLP or a related field, with some holding PhDs or professorial positions. Their ages ranged from 25 to 55.
The annotation guidelines were initially developed by the authors and subsequently refined in collaboration with the annotators to ensure a shared and well-defined set of evaluation criteria. Detailed instructions for the human annotators are provided in Fig. 3. The annotation template also included a comment section for uncertain cases, though no comments were submitted.
Appendix B Dataset Details
We include all *ACL conferences from 2017 to 2022 in NUTSHELL, covering main conferences, Findings, industry tracks, and workshops. As not all conferences are held every year, the number of talks varies accordingly. Table 5 provides a detailed overview.
Split Conference Year Talks
Train ACL 2017 140
2018 185
2019 244
2021 849
EMNLP 2017 93
2018 221
2021 1480
NAACL 2018 120
2019 114
2021 554
Dev ACL 2022 885
Test EMNLP 2022 465
NAACL 2022 966
Table 5: Number of talks per conferences in the NUTSHELL dataset.
Appendix C Baseline Details
@@
Appendix D Evaluation Details
@@
D.1 ROUGE and BERT Score
@@ D.2 LLM as a judge
Appendix E Human Evaluation for Model Outputs
We evaluate the models using ROUGE (Lin, 2004), BERTScore (Zhang et al., 2020), and LLM-as-a-judge. However, it is known that automatic evaluation metrics can come with limitations. Namely, the first two metrics may not fully capture semantic overlap (Deutsch and Roth, 2021), while LLM-as-a-judge is sensitive to prompt phrasing (Thakur et al., 2024) and struggles to distinguish between closely similar candidates (Shen et al., 2023). To validate the reliability of our automatic evaluation scores and better understand model behavior, we complement these metrics with a human evaluation. This allows us also to verify the robustness of our findings.
Specifically, we asked nine domain experts (four women and five men) to rank model outputs relative to the reference abstract, with each example annotated by three independent annotators. All annotators were already familiar with the NLP domain, scientific writing and presentation, and the task itself. They are experts in Natural Language Processing, holding at least a master’s degree in NLP or a related field, with some holding PhDs or professorial positions. Their ages ranged from 25 to 55. The annotation instructions are provided in Fig. 4.